OpenStudy (anonymous):

Hey! I need help Evaluating the integral of [5*(x^3)-7x+4]/[(x^2)*(x-1)] dx using partial fractions...

OpenStudy (tkhunny):

It's traditional to use division first to reduce the numerator to some degree less than the degree of the denominator.

OpenStudy (anonymous):

is this a long division problem, my original denominator was \[den=(x^{3}-x^{2})\] but that is linear, isn't it?

OpenStudy (tkhunny):

Both numerator and denominator are degree 3.

OpenStudy (anonymous):

wouldn't I go straight over to partial fractions?

OpenStudy (tkhunny):

Your life will be better if you resist that temptation. Division will give you at least one trivial term for your integrand. Why would you not want that?

OpenStudy (tkhunny):

Long division gives \(5 + \dfrac{5x^{2} - 7x + 4}{x^2(x-1)}\). See how nice and easy that '5' out front looks?!

OpenStudy (anonymous):

that 5 does look pretty nice...let me run through the long division to see if I have any problems on that

OpenStudy (tkhunny):

You sound converted. Sweet!!

OpenStudy (anonymous):

that was definitely the best response I've seen all week

OpenStudy (anonymous):

and I got the division correct, sweet!

OpenStudy (tkhunny):

Don't tell anyone, but we can actually have a little fun doing mathematics. Okay, now off to the Partial Fractions. What are your denominators?

OpenStudy (anonymous):

\[x, x^{2},(x-1)\] I believe that should be the decomposition

OpenStudy (tkhunny):

You have this. Pop it out and let's see what you get.

OpenStudy (anonymous):

Oh...that's why I flipped my biscuits...the way I would have done it would have had a cubic against a quadratic. I do believe I will get: \[\int\limits_{}^{}\frac{5x^2-7x+4}{x^3-x^2} dx=\int\limits_{}^{}\frac{3}{x}dx+\int\limits_{}^{}\frac{-4}{x^{2}}dx+\int\limits_{}^{}\frac{2}{(x-1)}dx\] which yields \[3\ln(\left|x\right|) + \frac{4}{x} + 2\ln(\left| x-1 \right|) + C\]

OpenStudy (tkhunny):

Except for one thing. Something else is entirely missing. What is it?

OpenStudy (anonymous):

+5 xD

OpenStudy (tkhunny):

That's it. We put it out there to be all pretty. No sense forgetting about it. Excellent work.

OpenStudy (anonymous):

I got a no-go from mathlab....are there any alternative forms

OpenStudy (tkhunny):

Perhaps I mistook your xD. That should be "5x".

OpenStudy (anonymous):

that deserves at least three awkard turtles. Thank you for the help!

OpenStudy (tkhunny):

Just a tiny glitch. Now we have it.